OpenStudy (anonymous):

Hey can you help me!! If u do i give your free medal!! Wat is the difference between 2 pi square r and pi square rsquared

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

write it out with the equation tool vvv down there

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

i needa get out my homework sheet

OpenStudy (anonymous):

cause i forgot

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

2pi r and pi r2 the difference

OpenStudy (anonymous):

In a circle with radius r, 2 pi r is the circumference, pi r (squared) is the area.

OpenStudy (agent0smith):

\[\Large C = 2 \pi r \]this gives circumference of a circle, C - the length around the outside \[\Large A=\pi r^2 \]This gives the area inside a circle.

OpenStudy (anonymous):

these? \[2 \pi r\] \[\pi r^2\] the first one is the circumference of a circle the second one is the area of a circle

OpenStudy (anonymous):

wat u mean by circumference?

OpenStudy (anonymous):

leme get pen out dfirst

OpenStudy (anonymous):

Circumference is the length of the line which forms the circle, as measured from one point once around the circle to the same point.

OpenStudy (anonymous):

ohh so the line that forms the circle?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

am i correct?

OpenStudy (anonymous):

The circumference measures the line around the circle's length.

OpenStudy (anonymous):

i still dont get it man can u explain in easy words because my english is bad!

OpenStudy (anonymous):

OK, I'll try. Take the circle and wrap a string around it. Mark the string once you're all the way around. Stretch the string out and measure between the marks. Does that help?

OpenStudy (anonymous):

OHHHHGH

OpenStudy (anonymous):

i get it thnxs!!

OpenStudy (anonymous):

btw

OpenStudy (anonymous):

can u draw a circle and list we alot of facts about circles? like draw them!

OpenStudy (anonymous):

??

OpenStudy (anonymous):

can you tewll me asap because yeh i have to go sleep nearly!@! And its due tmmr!

OpenStudy (agent0smith):

 Circumference is how long that line is. Area is how much space is inside the circle.

OpenStudy (anonymous):

ohh k u explain more easier for me but i already got it from Numb but thnxs!

OpenStudy (anonymous):